# Dome-DETR: DETR with Density-Oriented Feature-Query Manipulation for Efficient Tiny Object Detection

### v1

@@ 3.4. Progressive Adaptive Query Initialization
Existing DETR variants commonly employ a fixed number of queries (e.g., 300), selected based on token scores across the image. However, this static allocation fails to adapt to varying object densities, leading to inefficient computation in sparse regions and insufficient queries in dense areas. Previous works have attempted to address this issue by introducing categorized query numbers (Huang et al., 2024a) or dense queries (Zhang et al., 2023), yet they fail to dynamically adjust to object distributions, necessitating additional hyperparameter tuning and increasing adaptation complexity across datasets.
@@
We propose Progressive Adaptive Query Initialization (PAQI), a method that dynamically adjusts query density based on scene complexity while eliminating manual threshold tuning. Our approach builds on the observation that Top-K queries from the encoder output tend to be redundant in sparse regions while insufficient in dense areas. Additionally, many queries are initialized in background regions with no objects, contributing no benefit to detection. To address these inefficiencies, PAQI employs a progressive query initialization strategy that prioritizes high-response regions in the density map while filtering out low-response areas. This strategy eliminates dataset-specific hyperparameter tuning and enables adaptive query generation that aligns with the actual scene distribution.
@@
Candidate Query Selection via Classification Scores. As detailed in Algorithm 1, PAQI first generates a set of anchors based on the spatial distribution of the image. The encoded feature memory is processed through a classification head to obtain objectness scores, indicating the likelihood of each anchor containing an object. To balance redundancy reduction and adequate coverage in dense regions, we select the Top-$K_{M}$ features based on these scores as candidate queries.
@@
Query Splitting and Density-Aware Sampling. To enhance robustness across varying object densities, the selected queries are divided into two subsets: (1) a core set of the first $K_{N}$ queries, serving as a strong detection baseline, and (2) a flexible set containing the remaining $K_{M}-K_{N}$ queries for further refinement. We apply density-aware filtering using a high-response mask from DeFE, discarding low-response queries. This ensures that sparse regions are not oversampled while dense regions receive an appropriate query allocation.
@@
Bounding Box Generation and Dynamic NMS. To further regulate query numbers based on density, refined queries generate bounding box predictions through a regression head, followed by density-aware dynamic Non-Maximum Suppression (NMS).
@@
Traditional NMS uses a fixed IoU threshold, which can lead to excessive suppression in dense object regions or too many anchors in sparse areas. To mitigate this, we dynamically adjust the IoU threshold based on the classification score $\mathbf{S}_{final}$:
(14) $\mathbf{T}=IoU_{N}+\mathbf{S}_{final}\times(IoU_{M}-IoU_{N}),$
@@
where $\mathbf{S}_{final}$ is the score of the final mask-filtered tokens, and the IoU threshold is adapted based on object density. This prevents excessive suppression in dense areas while reducing query count in sparse regions. Specifically, a higher IoU threshold is applied in dense regions to mitigate over-suppression, whereas a stricter threshold is enforced in sparse regions to optimize computational efficiency, thereby enhancing detection performance and efficiency.
@@ 4.3. Implementation Details
Based on the D-FINE structure (Peng et al., 2024a), we employ a 1-layer transformer encoder, a deformable transformer decoder, and HGNetv2 as our CNN backbone (Zhao et al., 2024). Our model is trained on 8×8$\times$8 × NVIDIA 4090 GPUs. Following D-FINE (Peng et al., 2024a), we apply the same random crop and scale augmentation strategies. Additionally, we apply mixup (Zhang et al., 2017) and Mosaic (Bochkovskiy et al., 2020) techniques for VisDrone training, with Mosaic set to a probability of 1 and mixup set to a probability of 0.2.
@@ 4.4. Main Results
AI-TOD-V2. Table 2 shows our main results on the AI-TOD-V2 (Xu et al., 2022a) test split. We compare the performances of our Dome-DETR with strong baselines, including both non-end-to-end object detectors and end-to-end object detectors. All non-end-to-end object detectors except YOLOv12 use ResNet50 with feature pyramid network (FPN) (He et al., 2016). We re-implement a series of DETR-like models on AI-TOD-V2, and all DETR-like methods except DETR use 4-scale feature maps, which are extracted from stages 1, 2, 3, and 4 of the backbone. The results are summarized in Table 1, our proposed Dome-DETR-M and Dome-DETR-L achieves the best result 34.0 and 34.6 AP compared with other state-of-the-art methods, including non-end-to-end object detectors and end-to-end object detectors. Also, Dome-DETR-L surpasses the baseline by +3.7%, +4.6%, +1.4%, and +0.7% in terms of APvt, APt, APs, APm. The performance gain is greater on APvt, and APt, and our Dome-DETR outperforms the advanced series of DETR-like models on AI-TOD-V2. For LRP evaluation results, see Appendix A.
VisDrone. We also evaluate our Dome-DETR on the VisDrone dataset (Zhu et al., 2021). Table 2 presents results on the validation split, comparing Dome-DETR with state-of-the-art CNN-based and DETR-like detectors. Our Dome-DETR-L achieves the best AP (39.0%), surpassing D-FINE-L by +2.5% AP. Similarly, Dome-DETR-M outperforms D-FINE-M by +2.5% AP, while Dome-DETR-S exceeds D-FINE-S by +2.3% AP. Compared to DQ-DETR, Dome-DETR-L improves AP, AP50, and AP75 by +3.8%, +6.2%, and +5.2%, respectively.
@@ 4.6. Ablation Study
Hyperparameter Sensitivity Analysis. Table 4 highlights the impact of key hyperparameters on Dome-DETR-S performance. For IoUN and IoUM, the best AP (33.3) is achieved at (0.4, 0.9), as a moderate negative threshold prevents excessive suppression of true positives, while a stricter positive threshold ensures high-quality detections. Lowering IoUM to 0.7 slightly reduces AP due to increased ambiguity in positive assignments, whereas increasing IoUN to 0.6 marginally drops AP by over-filtering hard positives. For the query initialization threshold $T_{init}$, 0.05 provides the best balance, achieving AP 32.6. Raising it to 0.1 slightly decreases AP (32.9) due to excessive pruning, whereas lowering it to 0.03 significantly degrades AP (32.3) by introducing unnecessary background noise. For window size $H/W$, $H/W=10$ offers the best trade-off. A smaller $H/W$ enhances very tiny objects but lowers overall AP, while a larger $H/W$ benefits medium-sized objects at the cost of smaller object performance. This suggests an intermediate window resolution balances fine detail and global context.
@@ Appendix D Implementation Details
Number of attention layers in APE 1(×2)1($\times$ 2)1 ( × 2 ) 1(×2)1($\times$ 2)1 ( × 2 ) 1(×2)1($\times$ 2)1 ( × 2 )

### v2

@@ 3.4. Progressive Adaptive Query Initialization
Existing DETR variants typically use a fixed number of queries (e.g., 300), selected based on token scores. However, this static allocation fails to adapt to varying object densities, leading to inefficiency in sparse scenes and insufficient queries in dense ones. Prior works have attempted solutions via categorized query counts (Huang et al., 2024a) or dense queries (Zhang et al., 2023), but these lack adaptability to object distribution and require extra hyperparameter tuning, increasing complexity across datasets.
@@
We propose Progressive Adaptive Query Initialization (PAQI), which dynamically adjusts query density based on scene complexity without manual threshold tuning. Our method is based on the observation that Top-K encoder queries are often redundant in sparse regions and inadequate in dense ones. Many are also initialized in background regions, offering no benefit. To address this, PAQI progressively initializes queries by focusing on high-response regions in the density map while filtering out low-response areas. This eliminates dataset-specific tuning and enables adaptive, scene-aligned query generation.
@@
Candidate Query Selection via Classification Scores. As shown in Algorithm 1, PAQI first generates anchors based on the image’s spatial distribution. The encoded feature memory is passed through a classification head to obtain objectness scores, reflecting the likelihood of each anchor containing an object. To balance redundancy and coverage in dense areas, the Top-$K_{M}$ features by score are selected as candidate queries.
@@
Query Splitting and Density-Aware Sampling. To improve robustness under varying densities, the selected queries are split into: (1) a core set of the first $K_{N}$ queries as a strong baseline, and (2) a flexible set of the remaining $K_{M}-K_{N}$ for refinement. Density-aware filtering is applied using a high-response mask from DeFE, discarding low-response queries to avoid oversampling sparse regions while ensuring dense areas receive sufficient attention.
@@
Bounding Box Generation and Dynamic NMS. To further adapt query count to density, refined queries predict bounding boxes via a regression head, followed by density-aware dynamic Non-Maximum Suppression (NMS).
@@
Traditional NMS uses a fixed IoU threshold, which may cause excessive suppression in dense regions or too many anchors in sparse areas. To address this, we dynamically adjust the IoU threshold:
(14) $\mathbf{T}=IoU_{N}+\mathbf{D}\times(IoU_{M}-IoU_{N}),$
@@
where $\mathbf{D}$ is the predicted density of the final mask-filtered tokens, and the IoU threshold adapts to it. This avoids over-suppression in dense areas while reducing queries in sparse regions. Specifically, a higher IoU threshold is used in dense regions to prevent excessive suppression, and a stricter threshold in sparse regions to improve efficiency, enhancing both detection performance and speed.
@@ 4.3. Implementation Details
Based on the D-FINE structure (Peng et al., 2024a), we employ a 1-layer transformer encoder, a deformable transformer decoder, and HGNetv2 as our CNN backbone (Zhao et al., 2024). Our model is trained on $8\times$ NVIDIA 4090 GPUs. Following D-FINE (Peng et al., 2024a), we apply the same random crop and scale augmentation strategies. Additionally, we apply mixup (Zhang et al., 2017) and Mosaic (Bochkovskiy et al., 2020) techniques for VisDrone training, with Mosaic set to a probability of 1 and mixup set to a probability of 0.2.
@@ 4.4. Main Results
AI-TOD-V2. Table 2 presents our main results on the AI-TOD-V2 (Xu et al., 2022a) test split. We compare Dome-DETR with strong baselines, including both non-end-to-end and end-to-end detectors. Except for YOLOv12, all non-end-to-end detectors use ResNet50 with FPN (He et al., 2016). We re-implemented several DETR-like models on AI-TOD-V2; all except DETR use 4-scale feature maps from backbone stages 1–4. Results in Table 1 show our Dome-DETR-M and Dome-DETR-L achieve the best AP scores of 34.0 and 34.6, surpassing other state-of-the-art methods. Dome-DETR-L improves over the baseline by +3.7%, +4.6%, +1.4%, and +0.7% in $AP_{vt}$, $AP_{t}$, $AP_{s}$, and $AP_{m}$, with greater gains in $AP_{vt}$ and $AP_{t}$. Our model outperforms advanced DETR-like methods on AI-TOD-V2. LRP evaluation results are in Appendix A.
VisDrone. Table 2 presents results of VisDrone (Zhu et al., 2021) on the validation split. Our Dome-DETR-L achieves the best AP (39.0%), surpassing D-FINE-L by +2.5% AP. Similarly, Dome-DETR-M outperforms D-FINE-M by +2.5% AP, while Dome-DETR-S exceeds D-FINE-S by +2.3% AP. Compared to DQ-DETR, Dome-DETR-L improves AP, AP50, and AP75 by +3.8%, +6.2%, and +5.2%, respectively.
@@ 4.6. Ablation Study
Hyperparameter Sensitivity Analysis. Table 4 highlights the impact of key hyperparameters on Dome-DETR-S performance. For IoUN and IoUM, the best AP (33.3) is achieved at (0.4, 0.9), as a moderate negative threshold prevents excessive suppression of true positives, while a stricter positive threshold ensures high-quality detections. Lowering IoUM to 0.7 slightly reduces AP due to increased ambiguity in positive assignments, whereas increasing IoUN to 0.6 marginally drops AP by over-filtering hard positives. For the query initialization threshold $T_{\mathrm{init}}$, 0.05 provides the best balance, achieving AP 32.6. Raising it to 0.1 slightly decreases AP (32.9) due to excessive pruning, whereas lowering it to 0.03 significantly degrades AP (32.3) by introducing unnecessary background noise. For window size $H/W$, $H/W=10$ offers the best trade-off. A smaller $H/W$ enhances very tiny objects but lowers overall AP, while a larger $H/W$ benefits medium-sized objects at the cost of smaller object performance. This suggests an intermediate window resolution balances fine detail and global context.
@@ 5. Conclusion
Acknowledgements
This work was in part supported by the National Natural Science Foundation of China under grants 62472399, 62032006 and 62021001.
@@ Appendix D Implementation Details
Table 8. Additional comparisons with classic works on VisDrone validation split. Latency is measured on an NVIDIA T4 GPU unless otherwise noted. Results show our proposed Dome-DETR achieve comparable performance with much lower latency.
Method Input Size $\downarrow$ #Params. $\downarrow$ GFLOPs $\downarrow$ Latency (ms) $\downarrow$ AP (%) $\uparrow$ AP50 (%) $\uparrow$ AP75 (%) $\uparrow$
[…]
Dome-DETR-M (Ours) $\times800$ 24M 285 ¡15 36.1 59.8 39.4
Dome-DETR-L (Ours) $\times800$ 36M 376 ¡20 39.0 61.1 40.8
[…]
Appendix E Frequently Asked Questions During Review
E.1. Design Details of DeFE
The Density-Focal Extractor is a lightweight yet effective module designed to generate precise object density priors, which guide subsequent feature enhancement and query initialization. Its key design details and innovations are as follows:
Multi-scale dilated convolutions with strategic rates: DeFE employs three parallel dilated convolutions with rates of 1, 2, and 3. This design is purposeful: a dilation rate of 1 captures fine-grained local features (critical for tiny object boundaries), while rates of 2 and 3 expand the receptive field to aggregate contextual information from medium and larger ranges, respectively. By fusing these multi-scale features, DeFE effectively models both the detailed texture of tiny objects and their spatial distribution patterns in complex scenes (e.g., dense clusters or sparse arrangements).
End-to-end density probability generation: After feature fusion, a global average pooling layer aggregates high-level contextual cues, ensuring the density estimation is not biased by local noise. A 1×1 convolution followed by a sigmoid activation then compresses the fused features into a pixel-wise density probability map (range 0–1), where higher values indicate regions with a higher likelihood of containing objects. This map serves as a ”spatial attention guide” for subsequent modules, enabling the model to focus on foreground regions.
Efficiency optimization: Despite its multi-scale design, DeFE remains lightweight, adding only 0.8M parameters and 17.6 GFLOPs to the overall model. This efficiency is achieved through the use of depthwise separable convolutions (DS-conv) in the feature extraction stage, which drastically reduce computational overhead compared to standard convolutions, while bilinear upsampling preserves spatial resolution without excessive cost. This balance of performance and efficiency makes DeFE suitable for real-time applications.
E.2. Difference Between Previous Cluster-aware Cropping Based Methods
Some reviewers suggested comparing our method with classic approaches in the field of UAV object detection, such as Focus&Detect, YOLC, and UFPMP. These methods mainly rely on the Cluster-aware Cropping strategy: they first cluster and crop regions of interest in the image, then perform detection using detectors, and finally reorganize and map the detection boxes back to the original image. Although these methods have unique advantages, they have an upper limit in real-time performance. This is because their strategies introduce heavy CPU and I/O tasks, making it impossible to achieve good parallelism.
In contrast, Dome-DETR emphasizes using density maps to guide the enhancement of regional features and the progressive initialization of Object Queries, and completes object detection in the image in a single step, thereby ensuring high detection efficiency and low latency.
More comparisons with these methods are shown in Table 8.
E.3. Contribution of Each Component in PAQI
To clarify the specific contribution of each sub-module in the Progressive Adaptive Query Initialization (PAQI) framework, we conducted ablation experiments by incrementally adding components to the baseline model:
Candidate Query Selection: This component expands the initial query budget from 300 to 1500, significantly increasing the model’s capacity to capture potential object instances. This expansion directly boosts recall (AR improves from 46.0 to 49.1) by covering more spatial regions, especially in dense scenes where tiny objects are clustered. However, the larger query pool also introduces higher computational overhead, as more queries require additional feature interactions.
Query Splitting & Density-Aware Sampling: Building on the expanded query set, this module refines query allocation by distinguishing between ”core queries” (targeting high-density object regions) and ”flexible queries” (adapting to sparse areas). It suppresses redundant queries in background or low-information regions, reducing unnecessary computations (lower GFLOPs compared to the previous step). This refinement yields a notable AP improvement from 32.4 to 33.0, striking a better balance between coverage and efficiency.
Bounding Box Generation & Dynamic NMS: The final component introduces density-aware non-maximum suppression (NMS), which adjusts suppression thresholds based on local object density. In dense regions with overlapping boxes, it retains more true positives by relaxing the threshold, while in sparse regions, it tightens the threshold to filter false positives. This dynamic mechanism pushes AP to the highest value of 33.3, ensuring optimal accuracy-efficiency trade-offs, particularly in complex scenes with varying object distributions. distributions.
@@ E.3. Contribution of Each Component in PAQI
Number of attention layers in APE $1(\times 2)$ $1(\times 2)$ $1(\times 2)$